# ADA-DPM: A Neural Descriptors-based Adaptive Noise Filtering Strategy for SLAM

### v2

@@ Abstract
footnotetext: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
According to the main types of sensors used, mainstream SLAM algorithms can be divided into visual SLAM and LiDAR SLAM. Among them, visual SLAM usually relies on camera sensors. Benefiting from the advantages of cameras such as light weight, simple installation, and low cost, visual SLAM has been widely used in unmanned aerial vehicles (UAVs) and mobile robot platforms with limited load capacity[5, 6, 7]. However, the location accuracy of visual SLAM decreases significantly in situations with sudden changes in lighting conditions and lack of scene texture features. Compared with cameras[8, 9], the point cloud collected by LiDAR is less susceptible to environmental lighting conditions and can more effectively preserve the 3D structural information of the environment, making LiDAR SLAM one of the hot research directions attracting numerous scholars currently[10].
@@
Currently, according to different feature extraction methods, LiDAR SLAM can be mainly divided into handcrafted feature-based methods and deep learning-based methods. The former[11, 12, 13] relies on predefined geometric rules (such as corner features, plane features) to extract features, which easily leads to the problem of sparse or redundant features: in weak-texture scenes (such as tunnels, grasslands, etc.), geometric features are often sparse, which can easily cause the loss of key constraints and result in drift in pose estimation; while in scenarios with extremely rich geometric features (such as urban streets), although the features are dense, it is accompanied by huge computational overhead and high memory usage, making real-time reconstruction of large-scale scenes difficult[14]. In contrast, deep learning-based methods[15, 16] use neural networks to autonomously extract and fuse geometric features in the scene, enabling more compact and expressive feature representations, thus completing robot localization and mapping tasks more efficiently.
@@
However, in some complex scenarios (such as scenarios lacking fixed landmarks like buildings and roads; scenes with complex terrains like hills and gullies; and scenes with many dynamic objects), deep learning-based methods still face challenges. (1) Although deep learning-based methods are generally robust to a small number of dynamic objects (such as pedestrians and vehicles), when the proportion of dynamic objects in the scene is too high, their ego-motion will generate additional local constraints, leading to a decline in trajectory accuracy or even trajectory divergence. Existing dynamic point elimination methods (such as introducing auxiliary networks[17]) are effective, but inevitably increase the computational cost of the system. (2) In unstructured environments (such as hills and gullies), point cloud often contains a large amount of noise and outliers, which significantly affect the extraction of local features, cause incorrect matching of feature points, and ultimately lead to wrong pose estimation[18]. (3) In SLAM task, structurally feature points with regular distribution (such as those of buildings and road signs) are far easier to achieve correct matching than feature points with irregular distribution (for example, feature points of vegetation). Ignoring such differences will restrict the accuracy and efficiency of matching. (4) For the loop closure detection task in SLAM, using a single receptive field for feature extraction fails to fully utilize multi-scale neighborhood structural information. As a result, when the receptive field increases, the similarity of point clouds in certain overlapping regions may decrease significantly, which increases the probability of incorrect predictions for repeated scenes and ultimately leads to errors in pose estimation. In general, when existing methods deal with complex scenarios such as dynamic points, noise points, and unstructured environments, they often need to introduce additional constraints. This affects the matching of feature points to a certain extent, and they frequently face a trade-off between location accuracy and system robustness.
@@
To address the interference caused by dynamic objects, we propose the dynamic point prediction module (Dynamic Segmentation Head). It predicts the category of each feature point (whether it belongs to the dynamic point or static point) and directly removed the feature points predicted as dynamic points. This avoids dynamic points introducing incorrect local constraints, thereby eliminating the interference of the ego-motion of dynamic objects on the global pose prediction.
To address the interference caused by outliers and noise points, we propose the global importance scoring module (Importance Scoring Head). It assigns an importance score weight to each successfully matched feature point pair, emphasizing the influence of those reliable feature points that make significant contributions to point cloud registration, while filtering out low-importance points (which may be noise or outliers), thereby improving the accuracy of feature point matching and suppressing the negative impact of noise and unstructured feature points. Importantly, we propose the importance score loss, which consists of the global reconstruction loss weighted by importance score and the score loss, to supervise the prediction of global importance score weights. Firstly, the KAN Linear Layer[19] is used to directly predict the global rigid transformation of the point cloud to align feature points and compute the reconstruction error. Then, the predicted importance score weights are used to weight the reconstruction error, reducing the reconstruction error caused by low-importance points by predicting smaller importance score weights. Additionally, to prevent the model from reducing the global reconstruction loss by predicting all importance score weights as 0 during the training phase, we further construct the score loss with all labels set to 1 to force the model to predict non-zero score weights. Moreover, by supervising the prediction of global rigid transformations between feature points, implicit supervision is applied to the feature extraction network, encouraging the network to learn features that contribute more to the point cloud registration task.
@@ 3.1 Problem Definition
The core objective of LiDAR SLAM is to process the continuous frame point cloud data acquired by LiDAR, enabling real-time construction of an incremental environmental map and simultaneous estimation of the robot’s motion trajectory. The essence of this problem can be attributed to dynamic pose estimation and optimization: solving for the sensor’s motion trajectory through frame-to-frame point cloud registration in unknown environments. Its mathematical expression can be formalized as finding the optimal rigid transformation parameters, such that the current frame point cloud and the historical frame point cloud achieve spatial consistency after transformation.
Specifically, let the source point cloud of the $i$-th frame (current frame point cloud) be $\mathcal{P}_{src}=\{x_{s}\in\mathbb{R}^{3}\}_{s=1}^{S}$, and the target point cloud to be registered with it (usually selected from historical keyframe point clouds) be $\mathcal{P}_{dst}=\{y_{t}\in\mathbb{R}^{3}\}_{t=1}^{T}$ (where $x$ and $y$ are the 3D coordinates of points in $\mathcal{P}_{src}$ and $\mathcal{P}_{dst}$ respectively; $S$ and $T$ represent the number of points in the source point cloud and target point cloud respectively). The pose transformation $\mathcal{T}=\{\mathcal{R},t\}$ to be solved includes a rotation matrix $\mathcal{R}\in SO(3)$ and a translation vector $t\in\mathbb{R}^{3}$. The solution to the optimal transformation can be formulated as minimizing the following objective function:
$\min_{\mathcal{R},t}\sum_{x_{s},y_{t}\in\Psi}\left\|\mathcal{R}x_{s}+t-y_{t}\right\|_{2}$ (1)
where $\Psi$ represents the set of potential correspondences between $\mathcal{P}_{src}$ and $\mathcal{P}_{dst}$; $\|\cdot\|_{2}$ is the Euclidean distance.
However, in practical scenarios, point cloud data is inevitably disturbed by observation noise and dynamic object movement, making it difficult to directly obtain an accurate set of point correspondences $\Psi$. Moreover, the uncertainty of matched points causes traditional SLAM algorithms to easily fall into local optima when solving Equation (1). Therefore, the research focus of this paper is on how to avoid the interference of dynamic objects and noise, and establish accurate point correspondences $\Psi$. To facilitate the explanation of our method, we define the subscript ${\cdot_{src}}$ to represent feature descriptors from the source point cloud (i.e., the current frame point cloud), and ${\cdot_{dst}}$ to represent feature descriptors from the target point cloud (i.e., the historical frame point cloud); the symbol $\mathcal{P}$ represents three-dimensional coordinates; the symbol $\mathcal{F}$ represents features.
3.2 Overview of SLAM System
Figure 1: Overview of the SLAM system. The green part represents the odometry part; the blue part represents loop closure detection; the white part represents pose graph optimization; the blue-boxed portion within the white part represents the final global map and trajectory obtained.
@@ 3.2 Overview of SLAM System
As shown in Figure 1, the SLAM System mainly consists of three stages: Odometry, Loop-Closure Detection, and Pose Graph Optimization.
Odometry: The core task of the Odometry stage is to predict the optimal rigid transformation $\mathcal{T}_{i}$ of the current frame point cloud $\mathcal{P}_{i}$ (at the $i$-th moment, i.e. $\mathcal{P}_{src}$) relative to the historical frame point cloud set $\mathcal{P}$ (the historical frame point cloud set composed of the first $i-1$ frame point clouds, i.e., historical point clouds), so as to obtain the initial pose of $\mathcal{P}_{i}$. As shown in the green part of Figure 1, the inputs of Odometry include $\mathcal{P}_{i}$, $\mathcal{P}$, and $\mathcal{K}$ (the stored historical keyframe point cloud set). Firstly, we use KNN (K-Nearest Neighbor) to search in $\mathcal{P}$ for the point cloud closest to the keyframe point cloud $\mathcal{K}_{m}$ (i.e., the keyframe in the historical keyframe point cloud set $\mathcal{K}$ that is closest to $\mathcal{P}_{i}$), and denote it as the target point cloud $\mathcal{P}_{dst}$. Then, the point cloud registration algorithm is used to predict the rigid transformation $\mathcal{T}_{i}$ between $\mathcal{P}_{dst}$ and $\mathcal{P}_{i}$. Finally, the $\mathcal{T}_{i}$ and $\mathcal{P}_{i}$ are added to the pose graph for subsequent pose graph optimization.
@@
Meanwhile, during the Odometry, we also determine whether $\mathcal{P}_{i}$ should be added to $\mathcal{K}$ according to the following condition. Firstly, we use $\mathcal{T}_{i}$ to transform the successfully matched feature point pairs between $\mathcal{P}_{dst}$ and $\mathcal{P}_{i}$ into the same coordinate system, and calculate the RMSE (Root Mean Square Error, as shown in Equation (2)) between these matched feature point pairs; if the RMSE is less than a certain threshold and the matching confidence $\tau_{i}$ (for detailed definition, see Equations (3) and (4) in Section 3.C) of the two frame point clouds is greater than a certain threshold, then $\mathcal{P}_{i}$ is taken as a candidate keyframe point cloud. Secondly, to avoid the computational burden caused by excessively frequent insertion of keyframes, we calculate the minimum Euclidean distance of the poses between $\mathcal{P}_{i}$ and all keyframe point clouds in $\mathcal{K}$. Only when this distance is greater than a certain threshold, $\mathcal{P}_{i}$ is inserted into $\mathcal{K}$. It should be noted that in the initial stage of SLAM, the pose of the first frame point cloud is initialized as an identity matrix, and this frame point cloud is set as the first keyframe point cloud.
$RMES=\sqrt[]{\frac{1}{N_{p}}\sum\left[\mathcal{\hat{P}}_{dst}-\left(\mathcal{R}_{i}\mathcal{\hat{P}}_{i}+t_{i}\right)\right]^{2}}$ (2)
where, $N_{p}$ denotes the number of successfully matched feature points; $\hat{\mathcal{P}}_{dst}$ and $\hat{\mathcal{P}}_{i}$ represent the successfully matched feature points in $\mathcal{P}_{dst}$ and $\mathcal{P}_{i}$; $\mathcal{R}_{i}\in SO(3)$ and $t_{i}\in\mathbb{R}^{3}$ are the rotation and translation components of the rigid transformation $\mathcal{T}_{i}$.
Loop-Closure Detection: The Loop-Closure Detection stage detects and corrects the global trajectory drift caused by cumulative errors by predicting whether there exists a loop closure between $\mathcal{P}_{i}$ and $\mathcal{K}$. The specific process is shown in the blue part of Figure 1. Firstly, we calculate the Euclidean distance of the poses between $\mathcal{P}_{i}$ and $\mathcal{K}$, and find the set of loop closure candidate keyframes whose distance relative to $\mathcal{P}_{i}$ is less than a certain threshold. Then, we compute the relative pose between $\mathcal{P}_{i}$ and the candidate keyframe set, and exclude keyframes with small transformation magnitudes to obtain the loop closure keyframe set $\hat{\mathcal{K}}$. Next, we predict the probability of the loop closure between $\mathcal{P}_{i}$ and $\hat{\mathcal{K}}$, and select the top $k$ keyframe point clouds with the highest probability as the new loop closure keyframe set $\tilde{\mathcal{K}}$. Finally, we estimate the transformation pose between $\tilde{\mathcal{K}}$ and $\mathcal{P}_{i}$ using the point cloud registration algorithm, and update the predicted pose into the pose graph. It should be noted that to avoid frequent Loop-Closure Detection, Loop-Closure Detection will only be performed when the number of cumulatively inserted keyframes since the last Loop-Closure Detection reaches a certain threshold.
@@
Pose-Graph Optimization: The Pose-Graph Optimization stage (as shown in the white part of Figure 1) corrects the global trajectory drift by fusing historical observation data and utilizing global constraints. Firstly, using KNN, we search in the historical point cloud set $\mathcal{P}$ for the $k$ nearest point clouds to the most recent keyframe point cloud $\mathcal{K}_{m}$. By calculating their Euclidean distance to $\mathcal{P}_{i}$, we exclude point cloud frames that exceed a certain distance threshold, thus obtaining the set of neighboring frames of $\mathcal{P}_{i}$ in the pose graph; the point clouds in the neighboring frame set of $\mathcal{P}_{i}$ are transformed into the same coordinate system to obtain a local map. Then, we use the point cloud registration algorithm to compute the rigid transformation between $\mathcal{P}_{i}$ and the local map, which is used to optimize the pose in the pose graph. Finally, graph optimization is applied to optimize the pose graph, yielding the final rigid transformation pose (completed using the standard pose graph optimization provided in Open3D [37]).
@@ 3.3 ADA-DPM for Point Cloud Registration
Figure 2: Overview of the network architecture. The blue part represents the Encoder, and the white part represents the Decoder. The networks for both the source point cloud and the target point cloud share the same weights.
As shown in Figure 2, the ADA-DPM network mainly consists of two parts: Encoder and Decoder. The Encoder performs feature extraction on the input source point cloud $\mathcal{P}_{src}$ and target point cloud $\mathcal{P}_{dst}$, generating the sparse point cloud feature descriptor $[\overline{\mathcal{P}},\overline{\mathcal{F}}]$ containing compressed feature information. The Decoder predicts the rigid transformation between the two frames of point clouds and the probability of loop closure based on the feature descriptors $[\overline{\mathcal{P}},\overline{\mathcal{F}}]$.
Encoder: As shown in the blue part of Figure 2, the $\mathcal{P}_{src}$ and $\mathcal{P}_{dst}$ are fed into a weight-sharing feature extraction network, generating the sparse compressed point cloud descriptors for the source point cloud $[\overline{\mathcal{P}}_{src},\overline{\mathcal{F}}_{src}]$ and for the target point cloud $[\overline{\mathcal{P}}_{dst},\overline{\mathcal{F}}_{dst}]$ respectively. The feature extraction network consists of a weight-sharing PointNeXt [38] and a CLI-GCN module (see Section 3.D for details). PointNeXt downsamples the point cloud and aggregates adjacent geometric features, while the CLI-GCN module aggregates multi-scale local geometric features across receptive fields to enrich the structural information of the point cloud feature.
@@
Feature Attention Fusion Module: The feature attention fusion module consists of self-attention module, cross-attention module, and MLP. The self-attention module performs global feature integration on $[\overline{\mathcal{P}}_{src},\overline{\mathcal{F}}_{src}]$ and $[\overline{\mathcal{P}}_{dst},\overline{\mathcal{F}}_{dst}]$ encoded by the Encoder; the cross-attention module exchanges contextual information between the source point cloud and the target point cloud; the MLP integrates the feature information and outputs the decoded compressed point cloud feature descriptors $[\hat{\mathcal{P}}_{src},\hat{\mathcal{F}}_{src}]$ and $[\hat{\mathcal{P}}_{dst},\hat{\mathcal{F}}_{dst}]$.
@@
[…]
@@
where $\mathcal{M}_{s}$ represents the correlation matrix between $[\hat{\mathcal{P}}_{src},\hat{\mathcal{F}}_{src}]$ and $[\hat{\mathcal{P}}_{dst},\hat{\mathcal{F}}_{dst}]$; $\|\cdot\|_{r}$ and $\|\cdot\|_{c}$ denote row L2 normalization and column L2 normalization, respectively; @@@ represents matrix multiplication; $\mathcal{F}_{src}$ and $\mathcal{F}_{dst}$ are the features of the source and target point clouds obtained by concatenating $[\hat{\mathcal{P}}_{src},\hat{\mathcal{F}}_{src}]$ and $[\hat{\mathcal{P}}_{dst},\hat{\mathcal{F}}_{dst}]$ along the feature dimension, respectively; $\tau$ denotes the matching confidence between $[\hat{\mathcal{P}}_{src},\hat{\mathcal{F}}_{src}]$ and $[\hat{\mathcal{P}}_{dst},\hat{\mathcal{F}}_{dst}]$; $\operatorname{Softmax}_{r}(x)$ and $\operatorname{Softmax}_{c}(x)$ represent Softmax operations applied to rows and columns, respectively; $\operatorname{TopK}_{r}(x)$ denotes extracting the maximum value from each row; $\odot$ represents element-wise multiplication; $f_{\tau\rightarrow j}(\hat{\mathcal{P}}_{dst})$ is the mapping function that extracts the target point cloud compressed descriptor corresponding to $\tilde{\mathcal{P}}_{src}$ from $\hat{\mathcal{P}}_{dst}$ based on $\tau$.
@@
Dynamic Segmentation Head: The constructed Dynamic Segmentation Head consists of a multi-head attention module and an MLP & Sigmoid classifier, with its structure shown in Figure 3(a). The Dynamic Segmentation Head deeply mines the global information of $[\overline{\mathcal{P}},\overline{\mathcal{F}}]$ by simply stacking multi-head self-attention modules. After concatenation, the MLP & Sigmoid is used to predict the probability $\sigma_{d}\in\mathbb{R}^{N\times 2}$ that $[\overline{\mathcal{P}},\overline{\mathcal{F}}]$ belongs to dynamic points or static points (where $N$ represents the number of feature points in $[\overline{\mathcal{P}},\overline{\mathcal{F}}]$).

### v3

@@ 1 Introduction
According to the main types of sensors used, mainstream SLAM algorithms can be divided into visual SLAM and LiDAR SLAM. Among them, visual SLAM usually relies on camera sensors. Benefiting from the advantages of cameras such as light weight, simple installation, and low cost, visual SLAM has been widely used in unmanned aerial vehicles (UAVs) and mobile robot platforms with limited load capacity[5, 6, 7]. However, the location accuracy of visual SLAM decreases significantly in situations with sudden changes in lighting conditions and lack of scene texture features. Compared with cameras[8, 9, 10, 11, 12], the point cloud collected by LiDAR is less susceptible to environmental lighting conditions and can more effectively preserve the 3D structural information of the environment, making LiDAR SLAM one of the hot research directions attracting numerous scholars currently[13, 14].
@@
Currently, according to different feature extraction methods, LiDAR SLAM can be mainly divided into handcrafted feature-based methods and deep learning-based methods. The former[15, 16, 17, 18] relies on predefined geometric rules (such as corner features, plane features) to extract features, which easily leads to the problem of sparse or redundant features: in weak-texture scenes (such as tunnels, grasslands, etc.), geometric features are often sparse, which can easily cause the loss of key constraints and result in drift in pose estimation; while in scenarios with extremely rich geometric features (such as urban streets), although the features are dense, it is accompanied by huge computational overhead and high memory usage, making real-time reconstruction of large-scale scenes difficult[19]. In contrast, deep learning-based methods[20, 21] use neural networks to autonomously extract and fuse geometric features in the scene, enabling more compact and expressive feature representations, thus completing robot localization and mapping tasks more efficiently.
@@
However, in some complex scenarios (such as scenarios lacking fixed landmarks like buildings and roads; scenes with many dynamic objects), deep learning-based methods still face challenges. (1) In unstructured environments (such as rural and semi-urban scenes), point cloud often contains a large amount of noise and outliers, which significantly affect the extraction of local features, cause incorrect matching of feature points[22]. What’s more, structurally feature points with regular distribution (such as those of buildings and road signs) are far easier to achieve correct matching than feature points with irregular distribution (for example, feature points of vegetation). Ignoring such differences will restrict the accuracy and efficiency of matching. (2) Although deep learning-based methods are generally robust to a small number of dynamic objects (such as pedestrians and vehicles), when the proportion of dynamic objects in the scene is too high, their ego-motion will generate additional local constraints. Existing dynamic point elimination methods (such as introducing auxiliary networks[23]) are effective, but inevitably increase the computational cost of the system. (3) For the loop closure detection task in SLAM, using a single receptive field for feature extraction fails to fully utilize multi-scale neighborhood structural information. As a result, when the receptive field increases, the similarity of point clouds in certain overlapping regions may decrease significantly, which increases the probability of incorrect predictions for repeated scenes and ultimately leads to errors in pose estimation. In general, when existing methods deal with complex scenarios such as dynamic points, noise points, and unstructured environments, they often need to introduce additional constraints. This affects the matching of feature points to a certain extent, and they frequently face a trade-off between location accuracy and system robustness.
@@
To adaptively select feature points with higher contributions and suppress the influence of noise and unstructured feature points, the ADA-DPM introduces the Global Importance Scoring module. It assigns an importance score weight to each successfully matched feature point pair, emphasizing the influence of those reliable feature points that make significant contributions to point cloud registration, while filtering out low-importance points (which may be noise or outliers), thereby improving the accuracy of feature point matching and suppressing the negative impact of noise and unstructured feature points. Importantly, we propose the importance score loss, which consists of the global reconstruction loss weighted by importance score and the score loss, to supervise the prediction of global importance score weights. Firstly, the KAN Linear Layer[24] is used to directly predict the global rigid transformation of the point cloud to align feature points and compute the reconstruction error. Then, the predicted importance score weights are used to weight the reconstruction error, reducing the reconstruction error caused by low-importance points by predicting smaller importance score weights. Additionally, to prevent the model from reducing the global reconstruction loss by predicting all importance score weights as 0 during the training phase, we further construct the score loss with all labels set to 1 to force the model to predict non-zero score weights. Moreover, by supervising the prediction of global rigid transformations between feature points, implicit supervision is applied to the feature extraction network, encouraging the network to learn features that contribute more to the point cloud registration task.
To mitigate the interference caused by dynamic objects, we propose the dynamic point prediction module (Dynamic Segmentation Head). It predicts the category of each feature point (whether it belongs to the dynamic point or static point) and directly removed the feature points predicted as dynamic points. This avoids dynamic points introducing incorrect local constraints, thereby eliminating the interference of the ego-motion of dynamic objects on the global pose prediction.
@@ 2.2 Deep Learning-Based Methods
Figure 1: The SLAM pipeline of ADA-DPM. The green part represents the odometry part; the blue part represents loop closure detection; the white part represents pose graph optimization; the blue-boxed portion within the white part represents the final global map and trajectory obtained.
@@ 3.1 Problem Definition
The essence of LiDAR SLAM can be attributed to the optimization of dynamic pose estimation: solving for the robot’s motion trajectory through frame-to-frame point cloud registration in unknown environments. Its mathematical expression can be formalized as finding the optimal rigid transformation parameters $\mathcal{T}=\{\mathcal{R}\in SO(3),t\in\mathbb{R}^{3}\}$, such that the current frame point cloud and the historical frame point cloud achieve spatial consistency after transformation.
To facilitate the explanation of our method, we define the source point cloud (current frame point cloud) of the i-th frame as $\mathcal{P}_{src}=\{x_{s}\in\mathbb{R}^{3}\}_{s=1}^{S}$, and the target point cloud (typically selected from historical keyframe point clouds) to be registered with it as $\mathcal{P}_{dst}=\{y_{t}\in\mathbb{R}^{3}\}_{t=1}^{T}$ (where $x$ and $y$ are the 3D coordinates of points; $S$ and $T$ represent the number of points). We define the subscript $\cdot_{src}$} to represent feature descriptors from the source point cloud (i.e., the current frame point cloud), and $\cdot_{dst}$} to represent feature descriptors from the target point cloud (i.e., the historical frame point cloud); the symbol $\mathcal{P}$ represents three-dimensional coordinates; the symbol $\mathcal{F}$ represents features.
3.2 SLAM Pipeline of ADA-DPM
@@
As shown in Figure 1, SLAM System of ADA-DPM mainly consists of three stages: Odometry, Loop-Closure Detection, and Pose Graph Optimization.
@@
Odometry: The core task of the Odometry stage (as shown in the green part of Figure 1) is to predict the optimal rigid transformation $\mathcal{T}_{i}$ for the current frame point cloud $\mathcal{P}_{i}$ (at the $i$-th moment, i.e. $\mathcal{P}_{src}$) relative to the historical frame point cloud set $\mathcal{P}$ (the historical frame point cloud set composed of the first i1i-1 frame point clouds, i.e., historical point clouds), and to determine whether $\mathcal{P}_{i}$ should be added to the historical keyframe point cloud set $\mathcal{K}$. The specific process is as follows: first, the KNN (K-Nearest Neighbor) is used to search the set $\mathcal{K}$ for the point cloud frame closest $\mathcal{P}_{i-1}$, which is denoted as $\mathcal{P}_{dst}$. Then, the point cloud registration network in ADA-DPM is adopted to predict the $\mathcal{T}_{i}$ between $\mathcal{P}_{dst}$ and $\mathcal{P}_{i}$. Finally, based on the registration confidence $\tau_{i}$ (for detailed definition, see Equations (1) and (2) in Section 3.C) and RMSE (Root Mean Square Error) between $\mathcal{P}_{dst}$ and $\mathcal{P}_{i}$, it is determined whether to take $\mathcal{P}_{i}$ as a keyframe point cloud. In addition, to avoid the computational burden caused by excessively frequent keyframe insertion, $\mathcal{P}_{i}$ is only inserted into $\mathcal{K}$ when the minimum Euclidean distance of the pose between $\mathcal{P}_{i}$ and all frames in $\mathcal{K}$ is greater than a certain threshold.
Loop-Closure Detection: The core task of the Loop-Closure Detection stage (as shown in the blue part of Figure 1) is to construct a candidate loop frame set $\hat{\mathcal{K}}$ that may contain loops based on the Euclidean distance of poses between $\mathcal{P}_{i}$ and $\mathcal{K}$, and use the loop detection component in ADA-DPM to predict whether there exists a loop between $\mathcal{P}_{i}$ and $\hat{\mathcal{K}}$, so as to correct the global trajectory and map drift caused by accumulated errors. In addition, to avoid frequent loop-closure detection, our method only performs loop-closure detection when the number of keyframes accumulated since the last detection reaches a certain number. To minimize additional computation, keyframes with negligible pose differences are discarded during the candidate loop frame selection process.
@@
Pose-Graph Optimization: The Pose-Graph Optimization stage (as shown in the white part of Figure 1) fuses historical observation data and uses global constraints to correct trajectory and map drift. At first, we use the KNN to search the set $\mathcal{P}$ for the $k$ frames closest to $\mathcal{K}_{m}$. We then calculate the Euclidean distances between these $k$ frames and $\mathcal{P}_{i}$, discard the point cloud frames that exceed a specific distance threshold, and construct a local map composed of the neighboring frames of $\mathcal{P}_{i}$. Next, we use the point cloud registration network in ADA-DPM to compute the rigid transformation between $\mathcal{P}_{i}$ and the local map, and use this transformation to optimize the poses in the pose graph. Finally, we perform graph optimization on the pose graph to obtain the final global map (using the standard pose graph optimization provided in Open3D [44]).
@@ 3.3 ADA-DPM for Point Cloud Registration
Figure 2: The network overview of ADA-DPM. The blue part represents the Encoder, and the white part represents the Decoder. The networks for both the source point cloud and the target point cloud share the same weights.
As shown in Figure 2, the ADA-DPM network mainly consists of two parts: Encoder and Decoder. The Encoder performs feature extraction on the input source point cloud $\mathcal{P}_{src}$ and target point cloud $\mathcal{P}_{dst}$, generating the sparse point cloud feature descriptor $\overline{\mathcal{P}},\overline{\mathcal{F}}]$ containing compressed feature information. The Decoder predicts the rigid transformation between the two frames of point clouds and the probability of loop closure based on the feature descriptors $\overline{\mathcal{P}},\overline{\mathcal{F}}]$.
Encoder: As shown in the blue part of Figure 2, the $\mathcal{P}_{src}$ and $\mathcal{P}_{dst}$ are fed into a weight-sharing feature extraction network, generating the sparse compressed point cloud descriptors for the source point cloud $\overline{\mathcal{P}}_{src},\overline{\mathcal{F}}_{src}]$ and for the target point cloud $\overline{\mathcal{P}}_{dst},\overline{\mathcal{F}}_{dst}]$ respectively. The feature extraction network consists of a weight-sharing PointNeXt [45] and a CLI-GCN module (see Section 3.D for details). PointNeXt downsamples the point cloud and aggregates adjacent geometric features, while the CLI-GCN module aggregates multi-scale local geometric features across receptive fields to enrich the structural information of the point cloud feature.
@@
Feature Attention Fusion Module: The feature attention fusion module consists of self-attention module, cross-attention module, and MLP. The self-attention module performs global feature integration on $\overline{\mathcal{P}}_{src},\overline{\mathcal{F}}_{src}]$ and $\overline{\mathcal{P}}_{dst},\overline{\mathcal{F}}_{dst}]$ encoded by the Encoder; the cross-attention module exchanges contextual information between the source point cloud and the target point cloud; the MLP integrates the feature information and outputs the decoded compressed point cloud feature descriptors $\hat{\mathcal{P}}_{src},\hat{\mathcal{F}}_{src}]$ and $\hat{\mathcal{P}}_{dst},\hat{\mathcal{F}}_{dst}]$.
@@
In the point cloud registration process, the Dynamic Segmentation Head predicts and filters out dynamic points in $\overline{\mathcal{P}},\overline{\mathcal{F}}]$ output by the Encoder, before they enter the feature attention module. After the feature attention module, the Similarity Head performs feature point pair matching based on the similarity between $\hat{\mathcal{P}}_{src},\hat{\mathcal{F}}_{src}]$ and $\hat{\mathcal{P}}_{dst},\hat{\mathcal{F}}_{dst}]$, and predicts the matching confidence $\tau$ (mathematically described in Equations (1), (2)). The Offset Head predicts the offset between the compressed feature descriptors of the source point cloud and target point cloud after successful matching($\tilde{\mathcal{P}}_{src},\tilde{\mathcal{F}}_{src}]$ and $\tilde{\mathcal{P}}_{dst},\tilde{\mathcal{F}}_{dst}]$), eliminating the misalignment between them in the global coordinate system. Then, the Importance Scoring Head provides an initial estimation of the rigid transformation between $\tilde{\mathcal{P}}_{src},\tilde{\mathcal{F}}_{src}]$ and $\tilde{\mathcal{P}}_{dst},\tilde{\mathcal{F}}_{dst}]$, assigns contribution scores to the matched feature point pairs, and discards those with low contribution scores. Finally, the SVD is used to calculate the rigid transformation between the successfully matched point pairs.
@@
where $\mathcal{M}_{s}$ represents the correlation matrix between $\hat{\mathcal{P}}_{src},\hat{\mathcal{F}}_{src}]$ and $\hat{\mathcal{P}}_{dst},\hat{\mathcal{F}}_{dst}]$; $\|\cdot\|_{r}$ and $\|\cdot\|_{c}$ denote row L2 normalization and column L2 normalization, respectively; @@ represents matrix multiplication; $\mathcal{F}_{src}$ and $\mathcal{F}_{dst}$ are the features of the source and target point clouds obtained by concatenating $\hat{\mathcal{P}}_{src},\hat{\mathcal{F}}_{src}]$ and $\hat{\mathcal{P}}_{dst},\hat{\mathcal{F}}_{dst}]$ along the feature dimension, respectively; $\tau$ denotes the matching confidence between $\hat{\mathcal{P}}_{src},\hat{\mathcal{F}}_{src}]$ and $\hat{\mathcal{P}}_{dst},\hat{\mathcal{F}}_{dst}]$; $\operatorname{Softmax}_{r}(x)$ and $\operatorname{Softmax}_{c}(x)$ represent Softmax operations applied to rows and columns, respectively; $\operatorname{TopK}_{r}(x)$ denotes extracting the maximum value from each row; $\odot$ represents element-wise multiplication; $f_{\tau\rightarrow j}(\hat{\mathcal{P}}_{dst})$ is the mapping function that extracts the target point cloud compressed descriptor corresponding to $\tilde{\mathcal{P}}_{src}$ from $\hat{\mathcal{P}}_{dst}$ based on $\tau$.
@@
Dynamic Segmentation Head: The constructed Dynamic Segmentation Head consists of a multi-head attention module and an MLP & Sigmoid classifier, with its structure shown in Figure 3(a). The Dynamic Segmentation Head deeply mines the global information of $\overline{\mathcal{P}},\overline{\mathcal{F}}]$ by simply stacking multi-head self-attention modules. After concatenation, the MLP & Sigmoid is used to predict the probability $\sigma_{d}\in\mathbb{R}^{N\times 2}$ that $\overline{\mathcal{P}},\overline{\mathcal{F}}]$ belongs to dynamic points or static points (where $N$ represents the number of feature points in $\overline{\mathcal{P}},\overline{\mathcal{F}}]$).